OpenStudy (anonymous):

At a certain college, the probability that a student attends full-time is 0.320 and the probability that the student attends part-time is 0.680. At the same college, the probability that a student receives financial aid is 0.570 while the probability that a student does not receive financial aid is 0.430. If a student is randomly selected, what is the probability that the student attends full-time and does not receive financial aid?

OpenStudy (amistre64):

bayes ...

OpenStudy (anonymous):

.32 / .57

OpenStudy (amistre64):

.32(.43) maybe

OpenStudy (anonymous):

* my bad

OpenStudy (anonymous):

.32x.57

OpenStudy (anonymous):

.43 wrong word

OpenStudy (amistre64):

or its: \[\frac{.32(.43)}{.32(.43)+.68(.57)}\]

OpenStudy (luffingsails):

LOL

OpenStudy (luffingsails):

I'm going with P(A) * P(B) = 0.32 * 0.43

OpenStudy (amistre64):